# Do Black Holes With Generalized Entropy Violate Bekenstein Bound?

### v1

@@ I Introduction: Generalized Entropy and Bekenstein Bound
where, as per the common practice in literature, we have set $\hbar=c=k_{B}=1$. We will keep $G$ explicit since we will discuss modification of the gravitational constant later. In general relativity, we can easily check that a Schwarzschild black hole saturates this bound. The Bekenstein bound also holds in non-gravitational system (not surprising — note the absence of $G$ in the inequality), and is related to the positivity of the relative entropy [2, 3]. The quantum informational perspective of the Bekenstein bound is still an active area of research [4]. The generality of the bound suggests that it could be fundamental, at least in asymptotically flat spacetimes [^1] Its status in anti-de Sitter spacetimes are less clear. See [5].. On the other hand, in the literature, one finds many attempts to study what happens if we were to replace Bekenstein-Hawking entropy with various generalized entropies, such as Barrow entropy [6], Tsallis(-Cirto) entropy [7, 8], Kaniadakis entropy [9, 10, 11], and Sharma-Mittal entropy [12].
One of the motivations for contemplating these other possibilities is that the Bekenstein-Hawking entropy $S={A}/{4G}=4\pi G^{2}M^{2}$ is not extensive in its energy $M$ (that is, $S(\lambda M)\neq\lambda S(M)$), which means that its underlying statistics is probably not the Gibbs-Boltzmann entropy. As the logic goes, since the Bekenstein-Hawking entropy is not extensive anyway, one might as well consider other non-extensive forms that were first considered in non-gravitational systems, and use it to generalize the Bekenstein-Hawking form. (The Barrow case is somewhat special as it is motivated as a quantum gravity effect.) This certainly does not look like a strong motivation, and so we do not advocate for the applications of generalized entropy for black holes. While we need not believe in any of these extensions, it is important to study the effects of modifying Bekenstein-Hawking entropy on black hole physics. If the Bekenstein-Hawking entropy is in fact correct and therefore unique, then perhaps we can see what will go wrong by modifying it. (Similarly, modified gravity theories also often lead to subtle consistency issues.)
One thing that could potentially go very wrong is the Bekenstein bound. To see this, we consider for example, the Tsallis entropy, usually given in the context of black holes as [^2] This may not be the case, see, e.g., [13].
@@
$S_{T}^{1-\frac{1}{\delta}}\leqslant\left(\frac{A_{0}}{4G}\right)^{1-\frac{1}{\delta}}=\text{const.}$. (4)
Evidently, since for fixed $\delta$, the Tsallis entropy $S_{T}$ is an increasing function of $A$, we see that the bound fails for large black holes. This was first pointed out in [14] (the example we use here is similar to their computation for the Barrow case). In general then, as argued in [14, 15], when a generalized entropy is used in place of Bekenstein-Hawking one, we should not expect the Bekenstein bound to hold. Thus, if we take the position that the Bekenstein bound is fundamental, this would either imply that generalized entropies are not viable (and thus this can be used to constrain parameters of the theory; see e.g.,[16]), or that we cannot naively directly apply any generalized entropy with the usual assumption that $R=2GM$ and $E=M$. In fact, changing the entropy alone while keeping everything else unchanged has been long realized to be problematic and may be inconsistent (see, e.g., [17]; also see [18] for other criticisms in the context of cosmology), though this is still widely done in the literature.
@@
where $A$ is the area of the black hole horizon, and prime denotes derivative with respect to $A$. Since $G_{\text{eff}}$ is in general a function unless the horizon is fixed, it is a novel class of varying-$G$ theory. We shall refer to this approach as the GEVAG approach, short for “Generalized Entropy and VArying-$\mathbf{G}$ gravity”. In this work, because we only consider the Schwarzschild (static) solution, $G_{\text{eff}}$ is fixed and we need not worry about its variation. The Schwarzschild solution has event horizon located at $r=2G_{\text{eff}}M$.
@@ II Bekenstein Bound For Tsallis-Schwarzschild Black Hole
Therefore, Bekenstein bound holds if and only if
@@
This is an increasing function that reduces to $2\pi$ when $\delta\to 1/2$, consistent with the standard Bekenstein bound for Bekenstein-Hawking entropy. In fact, we can see from Fig.(1) that the original Bekenstein bound is satisfied provided $1/2<\delta\leqslant 1$. In the limit of large $\delta$, we have $C\to 4\pi$, twice the standard value.
@@
    Strong Bekenstein Bound: $0<C\leqslant 2\pi$.
@@ III Bekenstein Bound For Rényi-Schwarzschild Black Hole
The first law gives the thermodynamic energy:
@@
$a\coloneq\alpha GM^{2}.$ (25)
Since the arctan function tends to $\pi/2$ when its argument diverges, we see that this quantity tends to
$\lim_{a\to{1/16\pi}}E=\frac{\pi}{2}\lim_{a\to{1/16\pi}}\frac{1}{4\sqrt{\pi\alpha G}}=\frac{\pi M}{2}.$ (26)
@@
Again, we want to find the $C$ in the relaxed Bekenstein bound $S\leqslant CRE$. A direct computation yields
@@
As shown in Fig.(3), unlike the Tsallis case, this is a decreasing function, so it always satisfies $C\leqslant 2\pi$. Hence in this case the standard Bekenstein bound is strictly satisfied. In the limit $a\to 0$, we recover the standard value $2\pi$ for the Bekenstein constant.
@@ IV Conclusion: General Result
Let us now present the general result for any generalized entropy $S=f(A)/4G$. For this to satisfy $S\leqslant CRE$, with $R=2G_{\text{eff}}M$, it suffices to let the Bekenstein constant be
@@
This combination is clearly dimensionless and reduces to $2\pi$ in the GR limit. Furthermore, and crucially, we see that $C$ is bounded because $f(A)$ should be bounded by assumption (otherwise the generalized entropy $S=f(A)/4G$ diverges, which does not make physical sense for a finite system). Likewise, $G_{\text{eff}}=G/f^{\prime}(A)$ is nonzero as long as the entropy function $f(A)$ is differentiable and $f^{\prime}(A)$ is finite. This is sensible for the same reason: if $f^{\prime}(A)$ diverges, then $f$ is becoming unbounded, which is not physical. The energy $E$ should also be finite for any sensible thermodynamical system. Furthermore, since $E$ should recover $M$ in the GR limit, at least for small deviation away from the Bekenstein-Hawking entropy, $E$ is guaranteed to be finite for any generalized entropy. (If for any unlikely reason, $E$ is divergent but $S$ finite, then the Bekenstein bound is trivially satisfied.)

### v2

@@ I Introduction: Generalized Entropy and Bekenstein Bound
where, as per the common practice in literature, we have set $\hbar=c=k_{B}=1$. We will keep $G$ explicit since we will discuss modification of the gravitational constant later. In general relativity, we can easily check that a Schwarzschild black hole saturates this bound. The Bekenstein bound also holds in non-gravitational system (not surprising — note the absence of $G$ in the inequality), and is related to the positivity of the relative entropy [2, 3]. The quantum informational perspective of the Bekenstein bound is still an active area of research [4]. The generality of the bound suggests that it could be fundamental, at least in asymptotically flat spacetimes [^1] Its status in anti-de Sitter spacetimes are less clear. See [5].. On the other hand, in the literature, one finds many attempts to study what happens if we were to replace Bekenstein-Hawking entropy with various generalized entropies, such as Barrow entropy [6], Tsallis(-Cirto) entropy [7, 8], Kaniadakis entropy [9, 10, 11], and Sharma-Mittal entropy [12], as well as generalizations thereof [13].
One of the motivations for contemplating these other possibilities is that the Bekenstein-Hawking entropy $S={A}/{4G}=4\pi G^{2}M^{2}$ is not extensive in its energy $M$ (that is, $S(\lambda M)\neq\lambda S(M)$), which means that its underlying statistics is probably not the Gibbs-Boltzmann entropy. As the logic goes, since the Bekenstein-Hawking entropy is not extensive anyway, one might as well consider other non-extensive forms that were first considered in non-gravitational systems, and use it to generalize the Bekenstein-Hawking form. (The Barrow case is somewhat special as it was motivated as a quantum gravity effect.) This certainly does not look like a strong motivation, and so we do not advocate for the applications of generalized entropy for black holes. While we need not believe in any of these extensions, it is important to study the effects of modifying Bekenstein-Hawking entropy on black hole physics. If the Bekenstein-Hawking entropy is in fact correct and therefore unique, then perhaps we can see what will go wrong by modifying it. (Similarly, modified gravity theories also often lead to subtle consistency issues.)
One thing that could potentially go very wrong is the Bekenstein bound. To see this, we consider for example, the Tsallis entropy, usually given in the context of black holes as [^2] This may not be the case, see, e.g., [14, 15].
@@
$S_{T}^{1-\frac{1}{\delta}}\leqslant\left(\frac{A_{0}}{4G}\right)^{1-\frac{1}{\delta}}=\text{const.}$ (4)
Evidently, since for fixed $\delta$, the Tsallis entropy $S_{T}$ is an increasing function of $A$, we see that the bound fails for large black holes. This was first pointed out in [16] (the example we use here is similar to their computation for the Barrow case). In general then, as argued in [16, 17], when a generalized entropy is used in place of Bekenstein-Hawking one, we should not expect the Bekenstein bound to hold. Thus, if we take the position that the Bekenstein bound is fundamental, this would either imply that generalized entropies are not viable (and thus this can be used to constrain parameters of the theory; see e.g.,[18]), or that we cannot naively directly apply any generalized entropy with the usual assumption that $R=2GM$ and $E=M$. In fact, changing the entropy alone while keeping everything else unchanged has been long realized to be problematic and may be inconsistent (see, e.g., [20, 19]; also see [21] for other criticisms in the context of cosmology), though this is still widely done in the literature.
@@
where $A$ is the area of the black hole horizon, and prime denotes derivative with respect to $A$. Since $G_{\text{eff}}$ is in general a function unless the horizon is fixed, it is a novel class of varying-$G$ theory. We shall refer to this approach as the GEVAG approach, short for “Generalized Entropy and VArying-$\mathbf{G}$ gravity”. In this work, because we only consider the Schwarzschild (static) solution, $G_{\text{eff}}$ is fixed and we need not worry about its variation. Since the field equation of GEVAG has the same form as general relativity, the Schwarzschild solution has event horizon located at $r=2G_{\text{eff}}M$. For more details, we refer the readers to [22].
@@ II Bekenstein Bound For Tsallis-Schwarzschild Black Hole
Therefore, $S/RE\leqslant 2\pi$ holds if and only if
@@
Some clarifications are useful at this point. Let us first return to the original Bekenstein bound. The ratio $S/RE$ is strictly less than $2\pi$ for non-black hole systems, and is saturated by a black hole. In principle, $S/RE\leqslant C$ for any $C\geqslant 2\pi$ would also be mathematically correct, but physically $C>2\pi$ would correspond to a hypothetical “hyper-entropic” object (i.e. an object whose entropy exceeds a black hole of the same energy). Such an object is likely to collapse into a black hole [28, 29, 30, 31]. In fact, the Bekenstein bound means that such objects should be excluded — they would collapse into a black to restore the Bekenstein bound. For the same reason [^4] Admittedly there is a caveat: whether hyper-entropic objects are indeed unstable for the case of generalized entropy requires more detailed investigations., we define the “Bekenstein constant” $C$ as the value that is saturated by a black hole in the case with generalized entropy.
@@
Eq.(18) is an increasing function that reduces to $2\pi$ when $\delta\to 1$, consistent with the standard Bekenstein bound for Bekenstein-Hawking entropy. In fact, we can see from Fig.(1) that the original Bekenstein bound is satisfied for $\delta=1$, and that $C$ is strictly less than $2\pi$, indicating an even tighter bound than the original version, provided that $1/2<\delta\leqslant 1$. In the limit of large $\delta$, we have $C\to 4\pi$, twice the standard value.
@@
    Strong Bekenstein Bound: $0<C<2\pi$,
  • (2)
    (Original) Bekenstein Bound: $C=2\pi$,
@@ III Bekenstein Bound For Rényi-Schwarzschild Black Hole
Let us now derive the thermodynamic energy. The first law is $dE=TdS_{R}$. The Hawking temperature of the black hole is $T=1/8\pi G_{\text{eff}}M$ [22]. Substituting in the expression for $S_{R}$ in Eq.(23), we can use the chain rule to get
$dE=T\frac{dS_{R}}{dM}dM.$ (24)
Since
$\frac{dS_{R}}{dM}$ $=\frac{1}{\alpha}\frac{d}{dM}\left[\ln\left[1+\alpha\left(\frac{4\pi(2G_{\text{eff}}M)^{2}}{4G}\right)\right]\right]$ (25)
$=\frac{1}{\alpha}\left(\frac{1-\sqrt{1-16\pi\alpha GM^{2}}}{M\sqrt{1-16\pi\alpha GM^{2}}}\right),$
upon simplifying the expression we finally obtain
$dE=\frac{dM}{\sqrt{1-16\pi\alpha GM^{2}}}.$ (26)
Integrating this finally gives the thermodynamic energy:
@@
$a\coloneq\alpha GM^{2},$ (28)
so that the energy-mass relation can be written as
$E=M\left[\frac{1}{\sqrt{16\pi a}}\arctan\left(\frac{\sqrt{16\pi a}}{\sqrt{1-16\pi a}}\right)\right].$ (29)
We note that
$\lim_{a\to{1/16\pi}}\left[\arctan\left(\frac{\sqrt{16\pi a}}{\sqrt{1-16\pi a}}\right)\right]=\frac{\pi}{2},$ (30)
and
$\lim_{a\to{1/16\pi}}\frac{M}{\sqrt{16\pi a}}=M.$ (31)
Therefore
$\lim_{a\to{1/16\pi}}E=\frac{\pi M}{2}.$ (32)
@@
Again, we want to find the $C$ in the relaxed Bekenstein bound $S\leqslant CRE$. Since $C$ is defined to be the saturated value of the generalized Bekenstein bound, it can be obtained as $C=S/RE$. A direct computation yields
$C=\dfrac{\dfrac{1}{\alpha}\ln\left[1+\alpha\left(\dfrac{A}{4G}\right)\right]}{2G_{\text{eff}}M\left[\dfrac{1}{4\sqrt{\pi\alpha G}}\arctan\left(\dfrac{4\sqrt{\pi\alpha G}M}{\sqrt{1-16\pi\alpha GM^{2}}}\right)\right]}.$ (34)
Substituting in $A=4\pi(2G_{\text{eff}}M)^{2}$ and the expression for $G_{\text{eff}}$ in Eq.(21), we finally obtain in terms of $a$,
@@
As shown in Fig.(3), unlike the Tsallis case, this is a decreasing function, so it always satisfies $C\leqslant 2\pi$. In this case the strong Bekenstein bound is strictly satisfied. In the limit $a\to 0$, we recover the standard value $2\pi$ for the Bekenstein constant. It should be emphasized that it is crucial to make the distinction between the strong Bekenstein bound and the standard Bekenstein bound. In the latter, black holes saturate the bound, while other systems satisfy strict inequality. Whereas for the former, black holes saturate a bound whose value $C$ is less than $2\pi$. Due to this physical difference, it would be misleading to say that such generalized entropy black holes satisfy the original bound.
@@ IV Conclusion: General Result
Let us now present the general result for any generalized entropy $S=f(A)/4G$. For this to satisfy $S\leqslant CRE$, with $R=2G_{\text{eff}}M$ for the Schwarzschild black hole under GEVAG scheme [22], it suffices to let the Bekenstein constant be
@@
This combination is clearly dimensionless and reduces to $2\pi$ in the GR limit. Furthermore, and crucially, we see that $C$ is bounded because $f(A)$ should be bounded by assumption (otherwise the generalized entropy $S=f(A)/4G$ diverges, which does not make physical sense for a finite system). Likewise, $G_{\text{eff}}=G/f^{\prime}(A)$ is nonzero as long as the entropy function $f(A)$ is differentiable and $f^{\prime}(A)$ is finite. This is sensible for the same reason: if $f^{\prime}(A)$ diverges, then $f$ is becoming unbounded, which is not physical. The energy $E$ should also be finite for any sensible thermodynamical system. Furthermore, since $E$ should recover $M$ in the GR limit, at least for small deviation away from the Bekenstein-Hawking entropy, $E$ is guaranteed to be finite for any generalized entropy. (If for any unlikely reason, $E$ is divergent but $S$ finite, then the (weak) Bekenstein bound is trivially satisfied.)